# VMamba: Visual State Space Model

### v3

@@ Appendix C Details of the relationship between SS2D and Self-attention
$\mathbf{H}$ $\coloneqq\left[\mathbf{h_{a}};\ldots;\mathbf{h_{b}}\right]\in\mathbb{R}^{T\times D_{k}\times D_{v}},\>\text{where}\>\mathbf{h_{i}}\in\mathbb{R}^{D_{k}\times D_{v}}$ (14)
$\mathbf{Y}$ $\coloneqq\left[\mathbf{y_{a}};\ldots;\mathbf{y_{b}}\right]\in\mathbb{R}^{T\times D_{v}},\>\text{where}\>\mathbf{y_{i}}\in\mathbb{R}^{D_{v}}$ (15)
@@
where $\mathbf{M}\coloneqq\texttt{tril}(T,T)\in\{0,1\}^{T\times T}$ denotes the temporal mask matrix with the lower triangular portion of a $T\times T$ matrix set to 1 and elsewhere 0. It is evident that how matrices $\mathbf{Q}$, $\mathbf{K}$, and $\mathbf{V}$ are multiplied in Eq. 20 closely resembles the process in the self-attention module of Vision Transformers. Moreover, if $\mathbf{w}$ is in shape $(T,D_{k})$ rather than $(T,D_{k},D_{v})$, then Eq. 17 and Eq. 20 reduces to the form of Gated Linear Attention (GLA) [67], which indicates that GLA is also a special case of Mamba.
@@ Appendix D Visualization of Attention and Activation Maps
Given the input image as in Figure 8 (a), illustrations of four scanning paths in SS2D are shown in Figure 8 (d), the visualization of corresponding attention maps by calculating $\mathbf{Q}\mathbf{K}^{\top}$ and $\left(\mathbf{Q}\odot\mathbf{w}\right)\left(\mathbf{K}/\mathbf{w}\right)^{\top}$ are shown in Figure 8 (e) and Figure 8 (g) respectively. These results underscore the effectiveness of the proposed scanning approach (i.e., Cross-Scan) in capturing and retaining the traversed information, as each row in a single attention map corresponds to the attention between the current patch and all previously scanned foreground tokens impartially. Additionally, in Figure 8 (f), we showcase the transformed activation maps, where the pixel order corresponds to that of the first route, traversing the image row-wise from the upper-left to the bottom-right.
By rearranging the diagonal elements of the obtained attention map in the image space, we derive the visualization results illustrated in Figure 8 (b) and Figure 8 (c) corresponding to $\mathbf{Q}\mathbf{K}^{\top}$ and $\left(\mathbf{Q}\odot\mathbf{w}\right)\left(\mathbf{K}/\mathbf{w}\right)^{\top}$ respectively. These maps illustrate the effectiveness of VMamba in accurately distinguishing between foreground and background pixels within an image.
Moreover, given a certain selected patch as the query, we visualize the corresponding ‘activation map’ by reshaping the associated row in the attention map (computed by $\mathbf{Q}\mathbf{K}^{\top}$ or $\left(\mathbf{Q}\odot\mathbf{w}\right)\left(\mathbf{K}/\mathbf{w}\right)^{\top}$) This reflects the attention score between the query patch and all previously scanned patches. To get the complete visualization for a query patch, we collect and combine the activation maps from all four scanning paths in SS2D. Visualization results of the activation map for both $\mathbf{Q}\mathbf{K}^{\top}$ and $\left(\mathbf{Q}\odot\mathbf{w}\right)\left(\mathbf{K}/\mathbf{w}\right)^{\top}$ are shown in Figure 9. Moreover, we also visualize the diagonal elements of attention maps computed by $\left(\mathbf{Q}\odot\mathbf{w}\right)\left(\mathbf{K}/\mathbf{w}\right)^{\top}$, where all foreground objects are effectively highlighted and separated from the background.
@@ Appendix E Detailed Experiment Settings
The architectural specifications of Vanilla-VMamba are outlined in Table 3, and detailed configurations of the VMamba series are provided in Table 4. The Vanilla-VMamba series are built with the vanilla VSS Block, which includes a multiplicative branch with no feed-forward network (FFN) layer. In contrast, the VSS Block in the VMamba series eliminates the multiplicative branch and introduces FFN layers. Additionally, we provide alternative architectures for VMamba at Small and Base scales, referred to as VMamba-S[$s1l20$] and VMamba-B[$s1l20$], respectively. The notation ‘$sxly$’ indicates that the ssm-ratio is set to $x$ and the number of layers in stage 3 is set to $y$. Consequently, the versions presented in Table 1 can also be referred to as VMamba-S[$s2l15$] and VMamba-B[$s2l15$].
@@
The hyper-parameters for training VMamba on ImageNet are inherited from Swin [37], except for the parameters related to drop_path_rate and the exponential moving average (EMA) technique. Specifically, VMamba-T/S/B models are trained from scratch for 300 epochs, with a 20-epoch warm-up period, using a batch size of 1024. The training process utilizes the AdamW optimizer [39] with betas set to $(0.9,0.999)$, a momentum of $0.9$, an initial learning rate of $1\times 10^{-3}$, a weight decay of $0.05$, and a cosine decay learning rate scheduler. Additional techniques such as label smoothing (0.1) and EMA (decay ratio of 0.9999) are also applied. The drop_path_ratio is set to 0.2 for Vanilla-VMamba-T, VMamba-T, 0.3 for Vanilla-VMamba-S, VMamba-S[$s2l15$], VMamba-S[$s1l20$], 0.6 for Vanilla-VMamba-B, VMamba-B[$s2l15$] and 0.5 for VMamba-B[$s1l20$]. No other training techniques are employed.
@@
Detailed performance comparisons with various models are presented in Table 6. Throughput (referred to as TP.) was assessed on an A100 GPU paired with an AMD EPYC 7542 CPU, utilizing the toolkit provided by [64]. Following the protocol outlined in [37], we set the batch size to 128. The training throughput (referred to as Train TP.) is tested on the same device with mix-resolution, excluding the time consumption of optimizers. The batch size for measuring the training throughput is set to 128 as well.
@@
Table 5 provides detailed configurations of intermediate variants in the acceleration process from Vanilla-VMamba-T to VMamba-T.
@@
Table 3: Architectural overview of the Vanilla-VMamba series. Down-sampling is executed through patch merging [37] operations in stages 1, 2, and 3. The term Linear refers to a linear layer. The DWConv denotes a depth-wise convolution [24] operation. The proposed 2D-selective-scan is labeled as SS2D.
@@
Table 6: Performance comparison on ImageNet-1K with image size 224. indicates that Vim is trained only in float32 in practical, in which case the train throughput is 232. [71].
@@ Appendix F Performance of the VMamba Family on Downstream Tasks
For object detection and instance segmentation, we adhere to the protocol outlined by Swin [37] and construct our models using the mmdetection framework [3]. Specifically, we utilize the AdamW optimizer [39] and fine-tune the classification models pre-trained on ImageNet-1K for both 12 and 36 epochs. The learning rate is initialized with $1\times 10^{-4}$ and is decreased by a factor of 10 at the 9-th and 11-th epoch. We incorporate multi-scale training and random flipping with a batch size of 16, in line with established practices for object detection evaluations.
@@
Table 7: Object detection and instance segmentation results on COCO dataset. The FLOPs are calculated using inputs of size $1280\times 800$. Here, $AP^{b}$ and $AP^{m}$ denote box AP and mask AP, respectively. "1×1$\times$1 ×" indicates models fine-tuned for 12 epochs, while "3×3$\times$3 ×MS" signifies the utilization of multi-scale training for $36$ epochs.
@@
Table 8: Semantic segmentation results on ADE20K using UperNet [65]. We evaluate the performance of semantic segmentation on the ADE20K dataset with UperNet [65]. The FLOPs are calculated with input sizes of $512\times 2048$. "SS" and "MS" denote single-scale and multi-scale testing, respectively.
@@ Appendix G Details of VMamba’s Scale-Up Experiments
Given Mamba’s exceptional ability in efficient long sequence modeling, we conduct experiments to assess whether VMamba inherits this characteristic. We evaluate the computational efficiency and classification accuracy of VMamba with progressively larger input spatial resolutions. Specifically, following the protocol in XCiT [1], we apply VMamba trained on $224\times 224$ input to images with resolutions ranging from $288\times 288$ to $768\times 768$. We measure the generalization performance in terms of the number of parameters, FLOPs, throughput during both training and inference, and the top-1 classification accuracy on ImageNet-1K. We also conduct experiments under the ‘linear tuning’ setting, where only the header network, consisting of a single linear module, is fine-tuned from random initialization using features extracted by those backbone models.
According to the results summarized in Table 9, VMamba demonstrates the most stable performance across (i.e., modest performance drop) different input image sizes, achieving a top-1 classification accuracy of $74.7\%$ without fine-tuning ($79.2\%$ with linear tuning), while maintaining a relatively high throughput of 149 images per second at an input resolution of $768\times 768$. In comparison, Swin [37] achieves the second-highest performance with a top-1 accuracy of $73.1\%$ without fine-tuning ($77.5\%$ under linear tuning) at the same input size, when using scaled window sizes (set as the resolution divided by $32$). However, its throughput significantly drops to $53$ images per second. Furthermore, ConvNeXt [38] maintains a relatively high inference speed (i.e., a throughput of $103$ images per second) with the largest input resolution. However, its classification accuracy drops to $69.5\%$ when directly tested on images of size $768\times 768$, indicating its limited adaptability w.r.t. input images with large spatial resolutions. The performance of Deit-S also drops dramatically, mainly due to the use of interpolation in the absolute positional embedding.
The performance variations across different resolutions for various models are more intuitively illustrated in Figure 11. Notably, VMamba displays a linear increase in computational complexity (Figure 4 (b)), as measured by FLOPs, which is comparable to CNN-based architectures. This finding aligns with the theoretical conclusions drawn from selective SSMs [17].
@@
Figure 11: Throughtput and Top-1 accuracy change on ImageNet-1K w.r.t resolution rises. Swin-T denotes Swin-T tested with scaled window sizes.
@@ H.1 Influence of the Scanning Pattern
In the main submission, we validate the effectiveness of the proposed scanning pattern (referred to as Cross-Scan) in SS2D by comparing to three alternative image traversal approaches, i.e., Unidi-Scan, Bidi-Scan, and Cascade-Scan (Figure 12). Notably, as Unidi-Scan, Bidi-Scan, and Cross-Scan are all implemented in Triton, they exhibit little difference in throughput. Nevertheless, the findings in Table 10 indicate that Cross-Scan exhibits superior data modeling capacity, evident from its higher classification accuracy. This advantage is likely attributed to the two-dimensional prior introduced by the four-way scanning design. Nevertheless, the practical implementation of Cascade-Scan is notably constrained by its relatively slow computational pace. This limitation primarily stems from the inadequate compatibility between selective scanning and high-dimensional data, exacerbated by the multi-step scanning procedure.
Figure 13 indirectly demonstrates that among the analyzed scanning methods, only Bidi-Scan, Cascade-Scan, and Cross-Scan showcase global ERFs. Moreover, only Cross-Scan and Cascade-Scan exhibit two-dimensional (2D) priors. It is also noteworthy that DWConv [24] plays a significant role in establishing 2D priors, thereby contributing to the formation of global ERFs.
@@ H.2 Influence of the Initialization Approach
In our study, we adopted the initialization scheme initially proposed for the SS2D block in S4D [19]. Therefore, it is necessary to investigate the contribution of this initialization method to the effectiveness of VMamba. To delve deeper into this matter, we replaced the default initialization with two alternative methods: random initialization and zero initialization.
For both initialization methods, we set the parameter $\mathbf{D}$ in equation 1 to a vector of all ones, mimicking a basic skip connection (thus we have $\mathbf{y}=\mathbf{C}\mathbf{h}+\mathbf{D}\mathbf{u}$). Additionally, the weights and biases associated with the transformation from low rank to the dimension $D_{v}$ (which matches the input size), are initialized as random vectors, in contrast to Mamba [17], which employs a more sophisticated approach for initialization.
The main distinction between random and zero initialization lies in the parameter $A$ in equation 5, which is typically initialized as a HiPPO matrix in both Mamba [17, 19] and our implementation of VMamba. Given that we selected the hyper-parameter d_state to be 1, the Mamba initialization for $log(A)$ can be simplified to all zeros, which aligns with zero initialization. In contrast, random initialization assigns a random vector to $log(A)$. We choose to initialize $log(A)$ rather than $A$ directly to ensure that $A$ remains close to the all-ones matrix when the network parameters are near zero, which empirically helps enhance the stability of the training process.
The experiment results in Table 11 indicate that, at least for image classification with SS2D blocks, the model’s performance is not significantly affected by the initialization method. Therefore, within this context, the sophisticated initialization method employed in Mamba [17] can be substituted with a simpler, more straightforward approach. We also visualize the ERF maps of models trained with different initialization methods (see Figure 14), which intuitively reflect the robustness of SS2D across various initialization schemes.

### v4

@@ Appendix C Details of the relationship between SS2D and Self-attention
$\mathbf{H}$ $\coloneqq\left[\mathbf{h_{a+1}};\ldots;\mathbf{h_{b}}\right]\in\mathbb{R}^{T\times D_{k}\times D_{v}},\>\text{where}\>\mathbf{h_{i}}\in\mathbb{R}^{D_{k}\times D_{v}}$ (15)
$\mathbf{Y}$ $\coloneqq\left[\mathbf{y_{a+1}};\ldots;\mathbf{y_{b}}\right]\in\mathbb{R}^{T\times D_{v}},\>\text{where}\>\mathbf{y_{i}}\in\mathbb{R}^{D_{v}}$ (16)
@@
where $\mathbf{M}\coloneqq\texttt{tril}(T,T)\in\{0,1\}^{T\times T}$ denotes the temporal mask matrix with the lower triangular portion of a $T\times T$ matrix set to 1 and elsewhere 0. It is evident that how matrices $\mathbf{Q}$, $\mathbf{K}$, and $\mathbf{V}$ are multiplied in Eq. 21 closely resembles the process in the self-attention module of Vision Transformers. Moreover, if $\mathbf{w}$ is in shape $(T,D_{k})$ rather than $(T,D_{k},D_{v})$, then Eq. 18 and Eq. 21 reduce to the form of Gated Linear Attention (GLA) [65], indicating that GLA is also a special case of Mamba.
@@ Appendix D Visualization of Attention and Activation Maps
Given the input image shown in Figure 9 (a), illustrations of four scanning paths in SS2D are presented in Figure 9 (d). The visualizations of the corresponding attention maps, calculated using $\mathbf{Q}\mathbf{K}^{\top}$ and $\left(\mathbf{Q}\odot\mathbf{w}\right)\left(\mathbf{K}/\mathbf{w}\right)^{\top}$ are shown in Figure 9 (e) and Figure 9 (g) respectively. These results underscore the effectiveness of the proposed scanning approach (i.e., Cross-Scan) in capturing and retaining the traversed information, as each row in a single attention map corresponds to the attention between the current patch and all previously scanned foreground tokens impartially. Additionally, in Figure 9 (f), we showcase the transformed activation maps, where the pixel order corresponds to that of the first route, traversing the image row-wise from the upper-left to the bottom-right.
By rearranging the diagonal elements of the obtained attention map in the image space, we derive the visualization results shown in Figure 9 (b) and Figure 9 (c) corresponding to $\mathbf{Q}\mathbf{K}^{\top}$ and $\left(\mathbf{Q}\odot\mathbf{w}\right)\left(\mathbf{K}/\mathbf{w}\right)^{\top}$ respectively. These maps illustrate the effectiveness of VMamba in accurately distinguishing between foreground and background pixels within an image.
Moreover, given a selected patch as the query, we visualize the corresponding activation map by reshaping the associated row in the attention map (computed by $\mathbf{Q}\mathbf{K}^{\top}$ or $\left(\mathbf{Q}\odot\mathbf{w}\right)\left(\mathbf{K}/\mathbf{w}\right)^{\top}$) This reflects the attention score between the query patch and all previously scanned patches. To obtain the complete visualization for a query patch, we collect and combine the activation maps from all four scanning paths in SS2D. The visualization results of the activation map for both $\mathbf{Q}\mathbf{K}^{\top}$ and $\left(\mathbf{Q}\odot\mathbf{w}\right)\left(\mathbf{K}/\mathbf{w}\right)^{\top}$ are shown in Figure 10. We also visualize the diagonal elements of attention maps computed by $\left(\mathbf{Q}\odot\mathbf{w}\right)\left(\mathbf{K}/\mathbf{w}\right)^{\top}$, where all foreground objects are effectively highlighted and separated from the background.
@@ Appendix E Detailed Experiment Settings
The architectural specifications of Vanilla-VMamba are outlined in Table 3, while detailed configurations of the VMamba series are provided in Table 4. The Vanilla-VMamba series is constructed using the vanilla VSS Block, which includes a multiplicative branch and does not have feed-forward network (FFN) layers. In contrast, the VSS Block in the VMamba series removes the multiplicative branch and introduces FFN layers. Additionally, we provide alternative architectures for VMamba at Small and Base scales, referred to as VMamba-S[$s1l20$] and VMamba-B[$s1l20$], respectively. The notation ‘$sxly$’ indicates that the ssm-ratio is set to $x$ and the number of layers in stage 3 is set to $y$. Consequently, the versions presented in Table 1 can also be referred to as VMamba-S[$s2l15$] and VMamba-B[$s2l15$].
@@
The hyper-parameters for training VMamba on ImageNet are inherited from Swin [36], except for the parameters related to drop_path_rate and the exponential moving average (EMA) technique. Specifically, VMamba-T/S/B models are trained from scratch for 300 epochs, with a 20-epoch warm-up period, using a batch size of 1024. The training process utilizes the AdamW optimizer [38] with betas set to $(0.9,0.999)$, an initial learning rate of $1\times 10^{-3}$, a weight decay of $0.05$, and a cosine decay learning rate scheduler. It is noteworthy that this is not the optimal setting for VMamba. With a learning rate of $2\times 10^{-3}$, the Top-1 accuracy of VMamba-T can reach 80.7%.
Additional techniques such as label smoothing (0.1) and EMA (decay ratio of 0.9999) are also applied. The drop_path_ratio is set to 0.2 for Vanilla-VMamba-T and VMamba-T, 0.3 for Vanilla-VMamba-S, VMamba-S[$s2l15$] and VMamba-S[$s1l20$], 0.6 for Vanilla-VMamba-B and VMamba-B[$s2l15$], and 0.5 for VMamba-B[$s1l20$]. No additional training techniques are employed.
@@
Detailed performance comparisons with various models are presented in Table 6. Throughput (referred to as TP.) was assessed on an A100 GPU paired with an AMD EPYC 7542 CPU, utilizing the toolkit provided by [62]. Following the protocol outlined in [36], we set the batch size to 128. The training throughput (referred to as Train TP.) is tested on the same device with mix-resolution, excluding the time consumption of optimizers. The batch size for measuring the training throughput is also set to 128.
@@
Table 5 provides detailed configurations of the intermediate variants in the acceleration process from Vanilla-VMamba-T to VMamba-T.
@@
Table 3: Architectural overview of the Vanilla-VMamba series. Down-sampling is performed through patch merging [36] operations in stages 1, 2, and 3. The term Linear refers to a linear layer, while DWConv denotes a depth-wise convolution [23] operation. The proposed 2D-selective-scan is labeled as SS2D.
@@
Table 6: Performance comparison on ImageNet-1K with an image size of 224. indicates that Vim is trained solely in float32 in practice, with a training throughput of 232. [69].
@@ Appendix F Performance of the VMamba Family on Downstream Tasks
For object detection and instance segmentation, we adhere to the protocol outlined by Swin [36] and construct our models using the mmdetection framework [3]. Specifically, we utilize the AdamW optimizer [38] and fine-tune the classification models pre-trained on ImageNet-1K for both 12 and 36 epochs. The learning rate is initialized at $1\times 10^{-4}$ and decreased by a factor of 10 at the 9-th and 11-th epoch. We incorporate multi-scale training and random flipping with a batch size of 16, following established practices for object detection evaluations.
@@
Table 7: Object detection and instance segmentation results on COCO dataset. FLOPs are calculated using inputs of size $1280\times 800$. Here, $AP^{b}$ and $AP^{m}$ denote box AP and mask AP, respectively. "1×1$\times$1 ×" indicates models fine-tuned for 12 epochs, while "3×3$\times$3 ×MS" signifies the utilization of multi-scale training for $36$ epochs.
@@
Table 8: Semantic segmentation results on ADE20K using UperNet [63]. We evaluate the performance of semantic segmentation on the ADE20K dataset with UperNet [63]. FLOPs are calculated with input sizes of $512\times 2048$. "SS" and "MS" denote single-scale and multi-scale testing, respectively.
@@ Appendix G Details of VMamba’s Scale-Up Experiments
Given Mamba’s exceptional ability in efficient long sequence modeling, we conduct experiments to assess whether VMamba inherits this characteristic. We evaluate the computational efficiency and classification accuracy of VMamba with progressively larger input spatial resolutions. Specifically, following the protocol in XCiT [1], we apply VMamba, trained on $224\times 224$ inputs, to images with resolutions ranging from $288\times 288$ to $768\times 768$. We measure the generalization performance in terms of the number of parameters, FLOPs, throughput during both training and inference, and the top-1 classification accuracy on ImageNet-1K. We also conduct experiments under the ‘linear tuning’ setting, where only the header network, consisting of a single linear module, is fine-tuned from random initialization using features extracted by the backbone models.
According to the results summarized in Table 9, VMamba demonstrates the most stable performance across (i.e., modest performance drop) different input image sizes, achieving a top-1 classification accuracy of $74.7\%$ without fine-tuning ($79.2\%$ with linear tuning), while maintaining a relatively high throughput of 149 images per second at an input resolution of $768\times 768$. In comparison, Swin [36] achieves the second-highest performance with a top-1 accuracy of $73.1\%$ without fine-tuning ($77.5\%$ under linear tuning) at the same input size, using scaled window sizes (set as the resolution divided by $32$). However, its throughput significantly drops to $53$ images per second. Furthermore, ConvNeXt [37] maintains a relatively high inference speed (i.e., a throughput of $103$ images per second) at the largest input resolution. However, its classification accuracy drops to $69.5\%$ when directly tested on images of size $768\times 768$, indicating its limited adaptability to images with large spatial resolutions. Deit-S also shows a dramatic performance drop, primarily due to the interpolation used in the absolute positional embedding.
Notably, VMamba displays a linear increase in computational complexity, as measured by FLOPs, which is comparable to CNN-based architectures. This finding aligns with the theoretical conclusions drawn from selective SSMs [17].
@@ H.1 Influence of the Scanning Pattern
In the main submission, we validate the effectiveness of the proposed scanning pattern (referred to as Cross-Scan) in SS2D by comparing it to three alternative image traversal approaches, i.e., Unidi-Scan, Bidi-Scan, and Cascade-Scan (Figure 12). Notably, since Unidi-Scan, Bidi-Scan, and Cross-Scan are all implemented in Triton, they exhibit minimal differences in throughput. The results in Table 10 indicate that Cross-Scan demonstrates superior data modeling capacity, as reflected by its higher classification accuracy. This advantage likely stems from the two-dimensional prior introduced by the four-way scanning design. Nevertheless, the practical implementation of Cascade-Scan is significantly constrained by its relatively slow computational pace, primarily due to the inadequate compatibility between selective scanning and high-dimensional data, which is further affected by the multi-step scanning procedure.
Figure 13 indirectly demonstrates that among the analyzed scanning methods, only Bidi-Scan, Cascade-Scan, and Cross-Scan showcase global ERFs. Moreover, only Cross-Scan and Cascade-Scan exhibit two-dimensional (2D) priors. It is also worth noting that DWConv [23] plays a critical role in establishing 2D priors, thereby contributing to the formation of global ERFs.
@@ H.2 Influence of the Initialization Approach
In our study, we adopted the initialization scheme originally proposed for the SS2D block in S4D [19]. Therefore, it is necessary to investigate the contribution of this initialization method to the effectiveness of VMamba. To explore this further, we replaced the default initialization with two alternative methods: random initialization and zero initialization.
For both initialization methods, we set the parameter $\mathbf{D}$ in equation 1 to a vector of all ones, mimicking a basic skip connection (thus we have $\mathbf{y}=\mathbf{C}\mathbf{h}+\mathbf{D}\mathbf{u}$). Additionally, the weights and biases associated with the transformation to the dimension $D_{v}$ (which matches the input size), are initialized as random vectors. In contrast, Mamba [17] employs a more sophisticated initialization.
The main distinction between random and zero initialization lies in the parameter $A$ in equation 6, which is typically initialized as a HiPPO matrix in both Mamba [17, 19] and our implementation of VMamba. Given that we selected the hyper-parameter d_state to be 1, the Mamba initialization for $log(A)$ can be simplified to all zeros, which aligns with zero initialization. In contrast, random initialization assigns a random vector to $log(A)$. We choose to initialize $log(A)$ rather than $A$ directly to keep $A$ near the all-ones matrix when the network parameters are close to zero, which empirically enhances the training stability.
The experimental results in Table 11 indicate that, at least for image classification with SS2D blocks, the model’s performance is not significantly affected by the initialization method. Therefore, within this context, the sophisticated initialization method employed in Mamba [17] can be substituted with a simpler, more straightforward approach. We also visualize the ERF maps of models trained with different initialization methods (see Figure 14), which intuitively reflect SS2D’s robustness across various initialization schemes.